OpenStudy (anonymous):

lim as x approaches 0 of lnx-lnsinx

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (lnx-lnsinx)\]

OpenStudy (abb0t):

0

OpenStudy (anonymous):

\[\lim_{x\to0}\left(\ln x-\ln\sin x\right)=\lim_{x\to0}\ln\left(\frac{x}{\sin x}\right)\] I think you mean as x approaches 0 from the right... In any case, since \(\ln x\) is defined for \(x>0\), you can apply the continuity principle: \[\lim_{x\to0^+}\ln\left(\frac{x}{\sin x}\right)=\ln\left(\lim_{x\to0^+}\frac{x}{\sin x}\right)=\ln1\]

OpenStudy (anonymous):

Interesting, according to the graph, \(\ln\dfrac{x}{\sin x}\) *is* defined for \(x<0\)... Disregard that small note I made.